OpenStudy (anonymous):

The side length of a rectangular box with a square base is increasing at the rate of 2 ft/sec, while the height is decreasing at the rate of 2 ft/sec. At what rate is the volume of the box changing when the side length is 10 ft and the height is 12 ft? 680 ft3/sec 40 ft3/sec -280 ft3/sec 280 ft3/sec

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

If the base is fixed to be square as the side length increases, then the volume of the box is given by \[V=x^2z\] where \(x\) is the side length of the base and \(z\) is the height. Differentiating, you have \[\frac{dV}{dt}=2x\frac{dx}{dt}z+x^2\frac{dz}{dt}\] You're given that \(\dfrac{dx}{dt}=+2\) while \(\dfrac{dz}{dt}=-2\), and you want to find \(\dfrac{dV}{dt}\) at the time when \(x=10\) and \(z=12\).

OpenStudy (anonymous):

actually I meant to close this question because I figured it out, but if you can help with my other question that I posted today I would be super thankful!

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@SithsAndGiggles